# goal assembly as a formalism of evolvable design

### v1

@@ 1 Introduction
Consider what might be called the high-dimensional spherical cow model of phenotypic evolution, where variation is modeled as a high-dimensional random unit vector. In in $d>>1$ dimensions, two random vectors are almost orthogonal; the distribution of the component of the random variation vector that’s parallel to the fitness gradient limits a normal distribution with standard deviation $\sigma\sim 1/\sqrt{d}$. This scaling is easy to understand through symmetry arguments: there are $d-1$ orthogonal directions to the fitness gradient, but only one parallel. Then, upper bound the speed of directional evolution by selecting the single individual with the largest projection to the fitness gradient in each generation from a population size $n$. From extreme value statistics, we figure that this maximal projection scales as $\sigma\sqrt{\ln n}\sim\sqrt{\ln n}/\sqrt{d}$. This is the speed of evolution we can maximally get in the direction of the fitness gradient without any mechanism that facilitates variation to fall onto a lower-dimensional manifold. On the other hand, the projection to the direction that is orthogonal to the fitness gradient (i.e., neutral) is almost $1$. Of course, directional selection adds up over generations, and the neutral component is diffusive. Mathematically, with the number of generations $t$, evolution proceeds directionally (to the direction of the fitness gradient) as $t\sqrt{\ln n}/\sqrt{d}$, whereas in the neutral dimension as $\sqrt{t}$. It takes in the order of $d/\ln n$ generations for directional selection to take over. Mechanisms that meaningfully reduce the dimension of the search space help exponentially more than increasing population size. Arguments about high-dimensional phenotypic evolution, similar to this one, have a long history, starting with Fisher’s geometric model of adaptation [1, 2, 3, 4].
@@
How should such dimension reduction of phenotypic variation take place? Beyond varying the length of all four legs of a chair together, there is another, cognitive, principle. Consider the not so unreal example of an adversarial agent imposing complex perturbations on your LEGO or IKEA construction process. Soon you end up with a structure that’s not in the manual. Luckily, you find another manual that doesn’t encode a single construction trajectory, but all of them. It is very thick, and you end up spending hours finding a match to your structure, continuing your assembly process on that specific trajectory. Until, in five seconds, your adversary lands you on yet another new structure. The only reasonable strategy involves a paradigm shift in the creation of instruction manuals: encoding intermediate target structures, potentially multiple alternative ones that reflect the many possible parallel trajectories, and trusting the user’s common sense to reach them. Even better, users could share their own unique construction process, and with that feedback, intermediate goal states can start to align with what a human common sense can do, statistically. While current technology does not necessarily operate at a level of complexity and perturbations where this paradigm shift is necessary, this might not be the case in the future, especially if we actively embrace such a biologyfication of technology, in this abstract, substrate-independent sense.
@@ 2 Goal assembly formalism
There might be multiple non-conflicting explanations of this phenomena; I would like to point out how it might be, at least qualitatively, explained by dependency cascades in hierarchical systems. First, representational degeneracy, i.e., the number of local configurations that encode a given global configuration, is bounded by the number of non-coding local variables. For example, if a global configuration (phenotype), built from $N$ local variables overall, is encoded by only $C$ of the $N$ local variables, the degeneracy is at least $\prod_{i=1}^{N-C}m_{i}$ where $m_{i}$ are the number of possible states each non-coding local variable $i$, which is exponential in $N-C$. One way to have phenotypes with strongly varying degeneracies is to have a large enough variation in the number of non-coding local variables $N-C$. This can be achieved in hierarchical systems by dependency cascades: when a given goal variable is independent of at least one of the constituent goal variables, it is automatically independent of the full subtree of descendant goal variables, all the way down to the elementary goal variables. This has an exponential amplification effect: the higher in the hierarchy this independency happens, the larger the downstream effect is, resulting in more non-coding local variables.
@@
A simple probabilistic model of this amplification effect is the following. Assuming that the chance of independency at every step is an i.i.d. random variable, and also (wrongly) assuming that every dependency path the global variable to all local variables is independent, makes the distribution of the number of coding local variables $P(C)$ a product of i.i.d. variables, following a log-normal distribution. Another (analytically tractable) model could be formulated using coalescent theory, or stochastic branching processes, assuming that every goal variable depends on $k$ of their constituent goal variables where $k$ follows a Poisson distribution. In this case, the number of coding variables $P(C)$ at the local scale will also follow a Poisson distribution. But this is only the first step. The second step is to account for the fact that the number of non-coding variables map to the number of "neutral" configurations exponentially, as discussed above. This mapping is not totally trivial because we transform histograms, or probability density functions, so we have to follow their transformation rule. Consider the following back of the envelope calculation. Writing the typical number of goal states per variable $m$ is as $m=e^{\alpha}$, the size of local configuration space $s$ that represents a phenotype with $C$ coding local variables and $N-C$ non-coding ones is $s=e^{\alpha(N-C)}$, or equivalently, $C=N-\alpha^{-1}\ln s$. Then, from the transformation rule of probability density functions, $P(s)=P(C(s))|dC/ds|$, with $|dC/ds|=(\alpha s)^{-1}$. This, non-trivially, adds a "Zipfian" baseline to the representational bias distribution. For example, when $P(C)$ is uniform, $P(s)\sim s^{-1}$. Note that uniform $P(C)$ can be easily realized with a (near-)optimal encoding scheme, like prefix codes assigned by Huffman coding; however, since this is an optimization process involving the whole genotype space, it cannot be selected for, it must be emergent. What I offer here is a possible mechanistic explanation, using dependency cascades across the hierarchy, generating a wide variety of local coding variable numbers $C$ across phenotype realizations. This variety is, in turn, amplified when transformed to the share of genotype space s. Quantitatively, one might use the log-normal or Poisson model of $P(C)$ outlined above and transforms it according to $P(s)=\frac{P(C(s))}{\alpha s}$, with $C(s)=N-\alpha^{-1}\ln s$. For example, when we start from the narrower Poisson distribution, $P(C)\propto\frac{\lambda^{C}}{C!}$, we get (see Methods) $P(s)\propto\frac{s^{-\alpha^{-1}\ln\lambda-1}}{(N-\alpha^{-1}\ln s)!}$ which is a power law with a factorial cutoff, well approximated by the power law $P(s)\approx s^{-\alpha^{-1}\ln\lambda-1}$ when $s\lesssim e^{\alpha N}$, which is, approximately, the size of the whole genotype space. In other words, this "Zipfian baseline" seems to dominate the qualitative behavior of $P(s)$, but does not give us the numerical value of the exponent.
@@
Intuitively, compared to the configuration space, the "number of local variables" space is logarithmic, and the depth of the hierarchy is logarithmic in the number of local variables, and is double logarithmic in the number of configurations. This has a generic super-exponential or double-exponential configuration space reducing effect in the depth of the hierarchy. One simple example of that is shown by Fig 1f, where each goal variable selects $m$ of $m^{2}$ states as goal states, arranged in a binary tree, leading to a configuration space reduction of $m^{(2^{\ell})}$ in $l$ hierarchical integration (or goal composition) steps.
@@
Another core feature of goal assemblies is assumed to be a specific type of hierarchical modularity, in which components are often identical, as shown by Fig. 1e. "Vertical" encoding, using goal composition rules, as opposed to local configurations (Fig. 1a.), naturally accounts for this redundancy. This allows evolutionary innovation to happen in the space of compressed representations given by the goal integration rules. This kind of reliable encoding based innovation dynamics in the (open-ended) space of hierarchical modular structures is proposed to be a main focus of assembly theory [26], which is an additional reason why I call this current formalism goal assemblies, beyond the colloquial use of the word ’assembly’. Evolution in compressed representation space unites this ’pillar’ of goal assemblies with the previous one concerning statistics of genotype-to-phenotype or blueprint-to-construct maps (Fig. 1d.). Furthermore, joint decomposition of multiple phenotypes, as illustrated by Fig. 1e., can serve as a coordinate system, or map, which can be then used to formulate models of innovation dynamics in this compressed representation space. Historically contingent complexification dynamics ("in order to build something complex that works, you first need to build something simple that works"), with re-use and recombination of modules, here modules being minimal competencies, is a natural place where this framework might prove to be useful. As Fig. 2c. illustrates, evolvability in architecture space relies on a formalism that maps the ’adjacent possible’ of a population of hierarchical modular competency structures appropriately; goal assemblies provide a principled, compressed representation based approach.

### v2

@@ 1 Introduction
Consider what might be called the high-dimensional spherical cow model of phenotypic evolution, where variation is modeled as a high-dimensional random unit vector. In in $d>>1$ dimensions, two random vectors are almost orthogonal; the distribution of the component of the random variation vector that’s parallel to the fitness gradient limits a normal distribution with standard deviation $\sigma\sim 1/\sqrt{d}$. Then, upper bound the speed of directional evolution by selecting the single individual with the largest projection to the fitness gradient in each generation from a population size $n$. From extreme value statistics, we figure that this maximal projection scales as $\sigma\sqrt{\ln n}\sim\sqrt{\ln n}/\sqrt{d}$. This is the speed of evolution we can maximally get in the direction of the fitness gradient without any mechanism that facilitates variation to fall onto a lower-dimensional manifold. On the other hand, the projection to the direction that is orthogonal to the fitness gradient (i.e., neutral) is almost $1$. Of course, directional selection adds up over generations, and the neutral component is diffusive. Mathematically, with the number of generations $t$, evolution proceeds directionally (to the direction of the fitness gradient) as $t\sqrt{\ln n}/\sqrt{d}$, whereas in the neutral dimension as $\sqrt{t}$. It takes in the order of $d/\ln n$ generations for directional selection to take over. Mechanisms that meaningfully reduce the dimension of the search space help exponentially more than increasing population size. Arguments about high-dimensional phenotypic evolution, similar to this one, have a long history, starting with Fisher’s geometric model of adaptation [1, 2, 3, 4].
@@
How should such dimension reduction of phenotypic variation take place? Beyond varying the length of all four legs of a chair together, there is another, cognitive, principle. Consider the not so unreal example of an adversarial agent imposing complex perturbations on your LEGO or IKEA assembly process. Soon you end up with a structure that’s not in the manual. Luckily, you find another manual that doesn’t encode a single construction trajectory, but all of them. It is very thick, and you end up spending hours finding a match to your structure, continuing your assembly process on that specific trajectory. Until, in five seconds, your adversary lands you on yet another new structure. The only reasonable strategy involves a paradigm shift in the creation of instruction manuals: encoding intermediate target structures, potentially multiple alternative ones that reflect the many possible parallel trajectories, and trusting the user’s common sense to reach them. While current technology does not necessarily operate at a level of complexity and perturbations where this paradigm shift is necessary, this might not be the case in the future, especially if we actively embrace such a biologyfication of technology, in this abstract, substrate-independent sense.
@@ 2 Goal assembly formalism
There might be multiple non-conflicting explanations of this phenomena; I would like to point out how it might be, at least qualitatively, explained by dependency cascades in hierarchical systems. First, representational degeneracy, i.e., the number of local configurations that encode a given global configuration, is bounded by the number of non-coding local variables. For example, if a global configuration (phenotype), built from $N$ local variables overall, is encoded by only $C$ of the $N$ local variables, the degeneracy is at least $\prod_{i=1}^{N-C}m_{i}$ where $m_{i}$ are the number of possible states each non-coding local variable $i$, which is exponential in $N-C$. One way to have phenotypes with strongly varying degeneracies is to have a large enough variation in the number of non-coding local variables $N-C$. This can be achieved in hierarchical systems by dependency cascades: when a given goal variable is independent of at least one of the constituent goal variables, it is automatically independent of the full subtree of descendant goal variables, all the way down to the elementary goal variables. This has an exponential amplification effect: the higher in the hierarchy this independency happens, the larger the downstream effect is, resulting in more non-coding local variables. Writing the typical number of goal states per variable $m$ is as $m=e^{\alpha}$, the size of local configuration space $s$ that represents a phenotype with $C$ coding local variables and $N-C$ non-coding ones is $s=e^{\alpha(N-C)}$, or equivalently, $C=N-\alpha^{-1}\ln s$. Then, from the transformation rule of probability density functions, $P(s)=P(C(s))|dC/ds|$, with $|dC/ds|=(\alpha s)^{-1}$. This, non-trivially, adds a "Zipfian" baseline to the representational bias distribution. For example, when $P(C)$ is uniform, $P(s)\sim s^{-1}$. Note that uniform $P(C)$ can be easily realized with a (near-)optimal encoding scheme, like prefix codes assigned by Huffman coding; however, since this is an optimization process involving the whole genotype space, it cannot be selected for, it must be emergent. What I offer here is a possible mechanistic explanation, using dependency cascades across the hierarchy, generating a wide variety of local coding variable numbers $C$ across phenotype realizations. This variety is, in turn, amplified when transformed to the share of genotype space $s$. Consider a simple and conservative toy model, where the number of coding local variables $C$ is Poisson distributed, $P(C)\propto\frac{\lambda^{C}}{C!}$, with a standard deviation that scales as the square root of the mean $\lambda$. This can also be thought of as a lower bound on the width, generated by a non-cascading process, where the global phenotype depends on every local variable with a constant small probability. Any cascading process likely increases the width of this distribution. Since there is no reason to assume any further structure to the dependency cascade at this point, I provide an exact calculation of this lower bound here. Transforming $P(C)$ to $P(s)$ gives $P(s)\propto\frac{s^{-\alpha^{-1}\ln\lambda-1}}{(N-\alpha^{-1}\ln s)!}$ (see Methods) which is a power law with a factorial cutoff, well approximated by the power law $P(s)\approx s^{-\alpha^{-1}\ln\lambda-1}$ when $s\lesssim e^{\alpha N}$, which is, approximately, the size of the whole genotype space. In other words, this "Zipfian baseline" seems to dominate the qualitative behavior of $P(s)$, but does not give us the numerical value of the exponent.
@@
Intuitively, compared to the configuration space, the "number of local variables" space is logarithmic, and the depth of the hierarchy is logarithmic in the number of local variables, and is double logarithmic in the number of configurations. This has a generic super-exponential or double-exponential configuration space reducing effect in the depth of the hierarchy. One simple example of that is shown by Fig 1f, where each goal variable selects $m$ of $m^{2}$ states as goal states, arranged in a binary tree, leading to a configuration space reduction of $m^{(2^{\ell})}$ in $\ell$ hierarchical integration (or goal composition) steps.
@@
Another core feature of goal assemblies is assumed to be a specific type of hierarchical modularity, in which components are often identical, as shown by Fig. 1e. "Vertical" encoding, using goal composition rules, as opposed to local configurations (Fig. 1a.), naturally accounts for this redundancy. This allows evolutionary innovation to happen in the space of compressed representations given by the goal integration rules. This kind of reliable encoding based innovation dynamics in the (open-ended) space of hierarchical modular structures is proposed to be a main focus of assembly theory [26], which is an additional reason why I call this current formalism goal assemblies, beyond the colloquial use of the word ’assembly’. Evolution in compressed representation space unites this ’pillar’ of goal assemblies with the previous one concerning statistics of genotype-to-phenotype or blueprint-to-construct maps (Fig. 1d.). Furthermore, joint decomposition of multiple phenotypes, as illustrated by Fig. 1e., can serve as a coordinate system, or map, which can be then used to formulate models of innovation dynamics in this compressed representation space. Historically contingent complexification dynamics ("in order to build something complex that works, you first need to build something simple that works"), with re-use and recombination of modules, here modules being minimal competencies, is a natural place where this framework might prove to be useful. As Fig. 2d. illustrates, evolvability in architecture space relies on a formalism that maps the ’adjacent possible’ of a population of hierarchical modular competency structures appropriately; goal assemblies provide a principled, compressed representation based approach.